M (8,6) is the midpoint of RS. The coordinates of S are (9,7). What are the coordinates of R?: M (8,6) is the midpoint of RS. The coordinates of S are (9,7). What are the coordinates of R?
Answer by midwood_trail(310)

M (8,6) is the midpoint of RS. The coordinates of S are (9,7). What are the coordinates of R?
The midpoint formula is:
M = [(x1 + x2)/2, (y1 + y2)/2]
We are given the midpoint to be (8,6).
We are also given point S to be (9,7).
We set it up this way:
(8,6) = [(x + 9)/2, (y + 7)/2]
We now equate (x + 9)/2 to 8 and solve for x and then we equate (y + 7)/2 to 6 and solve for y.
(x + 9)/2 = 8
Multiply both sides by 2.
x + 9 = 8 (2)
x + 9 = 16
x = 16 - 9
x = 7
==========================
We now do the same for y.
(y + 7)/2 = 6
y + 7 = 6 (2)
y + 7 = 12
y = 12 - 7
y = 5
The coordinates of point R are (7, 5)
Did you follow?